A 100-volt electromotive force is applied to an RC-series circuit in which the resistance is 500 ohms and the capacitance is 10^−4 farad. Find the charge q(t) on the capacitor if q(0) = 0.

q(t) = _____

Find the current i(t).
i(t) = ____
1
Expert's answer
2020-06-24T18:22:14-0400

The Differential equation for RC-series circuit is

"R \frac{dq}{dt} + \frac{1}{C} q = E(t)" .

Now, "R = 500" ohm, C = "10^{-4\n}" F, E(t) = 100 V.

So, "500 \frac{dq}{dt} + 10^4 q = 100 \implies \frac{dq}{dt} + 20 q = 0.2"

This is standard first-order differential equation.

So, Integrating factor = "e^{\int20dt} = e^{20t}" .

Hence solution is "q e^{20t} = \int(0.2)e^{20t} dt + c"

"\implies qe^{20t} = (0.2) \frac{e^{20t} }{20}+ c =0.01 e^{20t} +c"

Now, "q(0)=0 \implies 0.01 + c = 0\implies c = -0.01"

"\implies qe^{20t} = 0.01(e^{20t}-1) \\ \implies q = \frac{0.01(e^{20t}-1)}{e^{20t}} = 0.01 (1-e^{-20t})"

And current "i(t) = \frac{dq}{dt} = 0.2 e^{-20t}"